OpenStudy (anonymous):

Write the following as a simple fraction in lowest terms. {(-6)^{-4} + (-6)^{-2}}/{-6^{-3}}

OpenStudy (aum):

\[\large \left((-6)^{-4} + (-6)^{-2}\right)/(-6)^{-3} \\ \text{To get rid of all the negative exponents, multiply top and bottom by } (-6)^4\\ \]

OpenStudy (anonymous):

ok why by (-6)^4

OpenStudy (aum):

The exponents are -2, -3, -4. If I multiply by the exponent +4, all three of the negative will become positive (or zero) because in a multiplication the exponents add. -2 + 4 = +2 -3 + 4 = +1 -4 + 4 = 0

OpenStudy (anonymous):

ok that makes sense

OpenStudy (aum):

\[\large \frac{(-6)^{-4} + (-6)^{-2}} {(-6)^{-3}} = \frac{(-6)^{-4} + (-6)^{-2}} {(-6)^{-3}}*\frac{(-6)^4}{(-6)^4} = \\ \large \frac{(-6)^{0} + (-6)^{2}} {(-6)^{1}} = \frac{1+36}{-6} = -\frac{37}{6} \\ \]

OpenStudy (anonymous):

ok thank you so much for your help that makes so much more sense

OpenStudy (aum):

You are welcome.